# Sublinear bilipschitz equivalence and the quasiisometric classification of solvable Lie groups

### v2

@@ 2.C. SBE Preserve Product Structure
$O(u)$-surjectivity. We now show that $\phi_{\hat{x}}$ is $v_{\hat{x}}-$quasi-surjective with $v_{\hat{x}}\in O(u)$. For convenience we show it in the case $\hat{x}=\hat{0}$, the base point of $\hat{X}=\prod_{m=2}^{n}X_{m}$. Assume towards contradiction that there is a sequence $y^{n}_{1}\in Y_{1}$ for which $\lim_{\omega}\frac{u(y^{n}_{1})}{d\big{(}\operatorname{Im}(\phi_{\hat{0}}),y^{n}_{1}\big{)}}=0$. Consider the points $y^{n}:=(y^{n}_{1},\hat{0}_{Y})$, where $\hat{0}_{Y}$ is the base point of $\hat{Y}:=\prod_{m=2}^{n}Y_{m}$. The quasi-surjectivity of $\phi$ implies that there are corresponding $x^{n}\in X$ for which $d\big{(}\phi(x^{n}),y^{n}\big{)}\leq u(y^{n})=u(y^{n}_{1})$. It is a general fact of sublinear bilipschitz embeddings that $x^{n}$ and $y^{n}_{1}$ admit $\frac{1}{L^{\prime}}|y^{n}_{1}|\leqslant|x^{n}|\leqslant L^{\prime}|y^{n}_{1}|$ for some $L^{\prime}>L$ and all large enough $y^{n}_{1}$.
Denote $\sigma_{n}:=d\big{(}\operatorname{Im}(\phi_{\hat{0}}),y^{n}_{1}\big{)}$, and consider the cones $X_{\omega}:=\operatorname{Cone}(X,x^{n},\sigma_{n})$, $Y_{\omega}:=\operatorname{Cone}\big{(}Y,\phi(x^{n}),\sigma_{n}\big{)}$. The previous paragraph implies $\lim_{\omega}\frac{u(x^{n})}{\sigma_{n}}=0$ and therefore by Lemma 2.7 the map $F=\operatorname{Cone}(\phi)$ is a bilipschitz homeomorphism. The hypothesis on the factors of $X$ and $Y$ implies that the map $F$ respects the product structure of $X_{\omega}$ and $Y_{\omega}$, i.e. there are bilipschitz maps $F_{i}$ for $1\leqslant i\leqslant n$ such that if $x_{\omega},x_{\omega}^{\prime}\in X_{\omega}$ agree on the $j$-coordinate, then
(2.3) $\pi_{(Y_{\omega})_{j}}\big{(}F(x_{\omega})\big{)}=F_{j}\big{(}\pi_{(X_{\omega})_{j}}(x_{\omega})\big{)}=F_{j}\big{(}\pi_{(X_{\omega})_{j}}(x_{\omega}^{\prime})\big{)}=\pi_{(Y_{\omega})_{j}}\big{(}F(x_{\omega}^{\prime})\big{)}$
Set $x_{\omega}:=(x^{n})_{n\in\mathbf{N}}$, $x^{0}_{\omega}:=(x^{n}_{1},\hat{0}_{\hat{X}})_{n\in\mathbf{N}}$ (where $x^{n}_{1}$ is the first coordinate of $x^{n}$). These two points in $X_{\omega}$ agree on the first coordinate, thus
$\Big{(}\pi_{Y_{1}}\big{(}\phi(x^{n}_{1},\hat{0})\big{)}\Big{)}_{n\in\mathbf{N}}=\pi_{(Y_{\omega})_{1}}\big{(}F(x_{\omega}^{0})\big{)}=\pi_{(Y_{\omega})_{1}}\big{(}F(x_{\omega})\big{)}=\pi_{(Y_{\omega})_{1}}\Big{(}\big{(}\phi(x^{n})\big{)}_{n\in\mathbf{N}}\Big{)}$
(Equalities are in $(Y_{\omega})_{1}$). On the other hand $d\big{(}\phi(x^{n}),(y^{n}_{1},\hat{0}_{Y})\big{)}\leqslant u(y^{n}_{1})$ and $\lim_{\omega}\frac{y^{n}_{1}}{\sigma_{n}}=0$, so $\big{(}\phi(x^{n})\big{)}_{n\in\mathbf{N}}=(y^{n}_{1},\hat{0}_{Y})_{n\in\mathbf{N}}$ (equality in $Y_{\omega}$). By definition $)\phi_{\hat{0}}(x^{n}_{1})=\pi_{Y_{1}}\big{(}\phi(x^{n}_{1},\hat{0}$$\big{)}$. We conclude
$\big{(}\phi_{\hat{0}}(x^{n}_{1})\big{)}_{n\in\mathbf{N}}=\pi_{(Y_{\omega})_{1}}\big{(}y^{n}_{1},\hat{0}_{Y})_{n\in\mathbf{N}}\big{)}=(y^{n}_{1})_{n\in\mathbf{N}},$
and from the Pythagorean formula we get $\lim_{\omega}\frac{d_{Y_{1}}\big{(}\phi_{\hat{0}}(x^{n}_{1}),y^{n}_{1}\big{)}}{\sigma_{n}}=0$, contradicting the definition of $\sigma_{n}$. We conclude there is a function $v_{\hat{0}}\in O(u)$ such that $\phi_{\hat{0}}$ is $v_{\hat{0}}-$quasi-surjective.
@@
$\big{(}\phi_{\hat{x^{n}}}(x^{n}_{1})\big{)}_{n\in\mathbf{N}}=\pi_{(Y_{\omega})_{1}}\big{(}F(x_{\omega})\big{)}=\pi_{(Y_{\omega})_{1}}\big{(}F(x_{\omega}^{0})\big{)}=\big{(}\phi_{\hat{0}}(x^{n}_{1})\big{)}_{n\in\mathbf{N}}$
And so $\lim_{\omega}\frac{d_{Y_{1}}\big{(}\phi_{\hat{x^{n}}}(x^{n}_{1}),\phi_{\hat{0}}(x^{n}_{1})\big{)}}{\sigma_{n}}=0$, contradicting the definition of $\sigma_{n}$. We conclude that indeed there is $v\in O(u)$ such that for all $\hat{x}\in\hat{X}$ and $x_{1}\in X_{1}$ we have
$d_{Y_{1}}\big{(}\phi_{\hat{x}}(x_{1}),\phi_{\hat{0}}(x_{1})\big{)}\leq v(|(\hat{x},x_{1})|).$
This inequality concludes the proof. ∎
As in [KKL98], in order to include Euclidean factors (or, more generally, factors with asymptotic cones homeomorphic to Euclidean space) we will also need a slight variation of the above result.
@@
Let $X:=\bar{X}\times Z,\ Y:=\bar{Y}\times W$ be two geodesic metric spaces. Assume that $f:X\rightarrow Y$ is an $(L,u)$-bilipschitz equivalence, and that for every pair of $u$-admissible cones $X_{\omega},Y_{\omega}$ the induced map $F=\operatorname{Cone}(f):X_{\omega}\rightarrow Y_{\omega}$ preserves the product $\bar{X}_{\omega}\times Z_{\omega}$ and $\bar{Y}_{\omega}\times W_{\omega}$ by the $Z_{\omega},W_{\omega}$ factors. Equivalently, there exists a homeomorphism $\bar{F}:\bar{X}_{\omega}\rightarrow\bar{Y}_{\omega}$ such that the following diagram commutes:
$X_{\omega}$$Y_{\omega}$$\bar{X}_{\omega}$$\bar{Y}_{\omega}$$\scriptstyle{F}$$\scriptstyle{\bar{\pi}_{\omega}}$$\scriptstyle{\bar{\pi}_{\omega}}$$\scriptstyle{\bar{F}}$
@@
Then there is an $(L^{\prime},v)$-bilispchitz equivalence $\bar{f}:\bar{X}\rightarrow\bar{Y}$, with $v\in O(u)$ and $L^{\prime}$ depending only on $L$ and $u$ such that the diagram
$X¯{\bar{X}}Y¯{\bar{Y}}$$\scriptstyle{f}$$\scriptstyle{\bar{\pi}}$$\scriptstyle{\bar{\pi}}$$\scriptstyle{\bar{f}}$
commutes up to sublinear $O(u)$-error.
Proof.
We define for each $z\in Z$ the map $\phi_{z}:\bar{X}\rightarrow\bar{Y}$ by $\phi_{z}(\bar{x})=\bar{\pi}(f(x,z))$. A similar computation to that done in the proof of Proposition 2.11 shows that these are $(L,u_{z})$-bilipschitz equivalences with $u_{z}\in O(u)$. To be a bit more explicit (but without repeating the entire argument): the quantity $d_{\bar{Y}}\big{(}\phi_{z}(\bar{x_{1}})\big{)},\phi_{z}(\bar{x_{2}})\big{)}$ is clearly bounded above by the upper bound on $d\big{(}f(\bar{x_{1}},z),f(\bar{x_{2}},z)\big{)}$ given by $f$. For the lower bound, the reverse triangle inequality implies that one has to find a sublinear function $v_{z}\in O(u)$ such that $d_{W}\Big{(}\pi_{W}\big{(}f(\bar{x_{1}},z)\big{)},\pi_{W}\big{(}f(\bar{x_{2}},z)\big{)}\Big{)}<v_{z}(\bar{x_{1}}\vee\bar{x_{2}})$. This is done in the usual way: a sequence of pairs $(\bar{x_{1}^{n}},z)$ and $(\bar{x_{2}^{n}},z)$ defying this property yields two points in an $u$-admissible cone which defy the fact that the cone map $F=\operatorname{Cone}(f)$ respects the product structure $X_{\omega}=\bar{X}_{\omega}\times Z_{\omega}$.
The quasi-surjectivity and sublinear control when varying $z$ are done exactly like in Proposition 2.11, concluding the proof. ∎
2.D. Concluding Theorem A
By assumption, the asymptotic cones of $X$ and $Y$ decompose into a product of a Euclidean factor and a non-Euclidean factor, where the non-Euclidean factor is a product of factors of types I and II. From [KKL98, Theorem 5.1] we see the cone maps between $u$-admissible cones preserve this product structure, and that the dimension of the Euclidean factors equal, i.e. $p=q$ in the notation of the statement. Proposition 2.18 gives rise to an $O(u)$-bilipschitz equivalence $\bar{\phi}:\prod_{i=1}^{n}X_{i}\rightarrow\prod_{j=1}^{m}Y_{j}$. Using [KKL98, Theorem 5.1] once more we see that $n=m$ and that in $u$-admissible cones the cone maps of $\bar{\phi}$ preserve the product structure. Theorem A now follows immediately from Proposition 2.11.
3. Completely solvable groups, Corollary B and Theorem C
3.A. Some preliminaries on distortion and Cornulier’s class
@@
    $H$ is a $O(\log)$-Lipschitz retract of $G$, more precisely: $\pi:G\to H$ is $O(\log)$-Lipschitz.
@@
We prove below that the groups in Example 1.7 are not quasiisometric. The proof is less involved than that of Theorem C but the main idea already intervenes, so we give it before.
@@
Let $\alpha\in(0,1)$. The groups $G^{0}_{4,9}$ and $\mathbf{R}\times G_{3,5}^{\alpha}$ are not quasiisometric, for any $\mu\neq 1$.
@@
Let’s check that $\mathbf{R}\times G_{3,3}$ is the group in the class $(\mathcal{C}_{1})$ associated to $G_{4,9}^{0}$ by [Cor11], and so there is a $O(\log)$-sublinear bilipschitz equivalence $\phi\colon G_{4,9}^{0}\to\mathbf{R}\times G_{3,3}$. The Lie algebra $\mathfrak{g}=\mathfrak{g}_{4,9}^{0}$ of $G_{4,9}^{0}$ is a semidirect product $\mathfrak{heis}\rtimes\delta$ where $\mathfrak{heis}$ has a basis $(e_{1},e_{2},e_{3})$ with $e_{1},e_{2}]=e_{3}$ and $\delta e_{i}=e_{i}$ if $i=1,3$ or 0 if $i=2$. The derived subalgebra is $\mathfrak{u}=\langle e_{1},e_{3}\rangle$ and $\mathfrak{g},\mathfrak{u}]=\mathfrak{u}$, so that $\mathfrak{u}$ is the Lie algebra of the exponential radical. The matrices of $\operatorname{ad}_{e_{3}}$ and $\operatorname{ad}_{e_{4}}$ in the basis $(e_{1},e_{2})$ are, respectively,
@@
From the definition of $\rho_{1}$ it follows that, using the same basis and changing the brackets, $\rho_{1}(\mathfrak{g}_{4,9})^{0}$ is the Lie algebra $\mathfrak{a}\rtimes\langle e_{4}\rangle\times\mathbf{R}\simeq\mathfrak{g}_{3,3}\times\mathbf{R}$.
Assume now towards contradiction that $G^{0}_{4,9}$ and $\mathbf{R}\times G_{3,5}^{\alpha}$ are quasiisometric; then by Cornulier’s Theorem 3.7 there is a $O(\log)$-bilispchitz equivalence $\psi\colon\mathbf{R}\times G_{3,3}\to\mathbf{R}\times G_{3,5}^{\alpha}$. By Theorem A there is a sublinear bilipschitz equivalence $\underline{\psi}\colon G_{3,3}\to G_{3,5}^{\alpha}$. However, by [Pal20b], $O(u)$-sublinear bilipschitz equivalences between Heintze groups preserve the conformal dimensions of their Gromov boundaries. $G_{3,3}$ and $G_{3,5}^{\alpha}$ are Heintze groups, and the conformal dimensions are
@@
3.C. Proof of Corollary B and Theorem C

### v3

@@ 2.C. Proof of Theorem C
$O(u)$-surjectivity. We claim that there is $v\in O(u)\textsterling$ such that for all $\hat{x}\in\widehat{X},y^{1}\in Y_{1}$, it holds that $d_{Y_{1}}(\operatorname{Im}\phi_{\hat{x}},y)\leq v(|\hat{x}|\vee|y^{1}|)$. Assume towards contradiction that this is not the case. This means there are sequences $\widehat{x_{k}}\in\widehat{X},y_{k}^{1}\in Y_{1}$ for which $\lim\frac{u(|\widehat{x_{k}}|\vee|y_{k}^{1}|)}{d_{Y_{1}}(\operatorname{Im}\phi_{\hat{x}},y_{k}^{1})}=0$.
We make the following construction and notation:
  • Let $x_{k}^{1}\in X_{1}$ be points realizing $d_{Y_{1}}\big(\phi_{\widehat{x_{k}}}(x_{k}^{1}),y_{k}^{1}\big)\leq d_{Y_{1}}(\operatorname{Im}\phi_{\widehat{x_{k}}},y_{k}^{1})+1$.
  • $\sigma_{k}:=d_{Y_{1}}\big(\phi_{\widehat{x_{k}}}(x_{k}^{1}),y_{k}^{1}\big)$.
@@
  • $\widehat{y_{k}}:=\pi_{\widehat{Y}}\circ\phi(\widehat{x_{k}},x_{k}^{1})$.
[…]
$\frac{1}{L}|x_{k}^{1}|-g(|x_{k}^{1}|)-g(|\widehat{x_{k}}|)\leq d_{Y_{1}}\big(\phi_{\widehat{x_{k}}}(x_{k}^{1}),\phi_{\widehat{x_{k}}}(0_{X_{1}})\big)\leq d_{Y_{1}}\big(\phi_{\widehat{x_{k}}}(x_{k}^{1}),0_{Y_{1}})\big)+d_{Y_{1}}\big(0_{Y_{1}},\phi_{\widehat{x_{k}}}(0_{X_{1}})\big).$
[…]
We also have the following estimates:
[…]
By our assumption that $\lim_{k}\frac{u(|\widehat{x_{k}}|\vee|y_{k}^{1}|)}{\sigma_{k}}=0$, we therefore have:
(2.4) $\lim_{k}\frac{u(|x_{k}|)+u(|z_{k}|)+u(|y_{k}|)}{\sigma_{k}}=0.$
Moreover, using the defining inequality of $\phi$ and the formula for the distance in product spaces, we have:
(2.5) $\begin{split}\frac{d_{X}(x_{k},z_{k})}{L}-u(|z_{k}|\vee|x_{k}|)&\leq d_{Y}\big(\phi(x_{k}),\phi(z_{k})\big)\\ &\lesssim d_{Y_{1}}\big(\phi_{\widehat{0}}(x_{k}^{1}),\pi_{Y_{1}}\circ\phi(z_{k})\big)+d_{\widehat{Y}}\big(\pi_{\widehat{Y}}\circ\phi(x_{k}),\pi_{\widehat{Y}}\circ\phi(z_{k})\big)\\ &\leq d_{Y_{1}}\big(\phi_{\widehat{0}}(x_{k}^{1}),y_{k}^{1}\big)+d_{Y_{1}}\big(y_{k}^{1},\pi_{Y_{1}}\circ\phi(z_{k})\big)+d_{\widehat{Y}}\big(\widehat{y_{k}},\pi_{\widehat{Y}}\circ\phi(z_{k})\big)\\ &\leq\sigma_{k}+u(y_{k})+u(y_{k})\lesssim\sigma_{k},\end{split}$
where the last inequality is a consequence of our assumption that $\lim\frac{u(|y_{k}^{1}|)}{\sigma_{k}}=0$.
Combining (2.5) with the estimates in (2.3) we conclude:
(2.6) $d_{X}(x_{k},z_{k})\lesssim\sigma_{k}+g(|z_{k}|\vee|x_{k}|)\lesssim\sigma_{k}+g(|y_{k}^{1}|)+g(|\widehat{x_{k}}|)+g(\sigma_{k})\lesssim\sigma_{k},$
where again the last inequality follows from (2.4) and $g\in O(u)$. Consider the cone $X_{\omega}:=\operatorname{Cone}(X,x_{k},\sigma_{k})$. Inequality (2.4) establishes that $X_{\omega}$ is $u$-admissible. Moreover, (2.6) ensures that the points $z_{\omega}:=[(z_{k})_{k}]$ and $z_{\omega}^{x}:=[(z_{k}^{1},\widehat{x_{k}})_{k}]$ both lie in $X_{\omega}$. Therefore by Lemma 2.15 we get
$\lim_{\omega}\frac{1}{\sigma_{k}}d_{Y_{1}}\big(\pi_{Y_{1}}\circ\phi(z_{k}^{1},\widehat{x_{k}}),\pi_{Y_{1}}\circ\phi(z_{k}^{1},\widehat{z_{k}})\big)=0,$
and therefore
$\lim_{\omega}\frac{1}{\sigma_{k}}d_{Y_{1}}\big(\phi_{\widehat{x_{k}}}(z_{k}^{1}),y_{k}^{1}\big)=\lim_{\omega}\frac{1}{\sigma_{k}}d_{Y_{1}}\big(\pi_{Y_{1}}\circ\phi(z_{k}),y_{k}^{1}\big)\leq\lim\frac{u(|y_{k}|)}{\sigma_{k}}=0$
(again, the last step is by Inequality (2.4)). We reach a contradiction to the definition of $\sigma_{k}<d(\operatorname{Im}\phi_{\widehat{x_{k}}},y_{k}^{1})+1$. We conclude the existence of $v\in O(u)$ as claimed.
Sublinear Control. We show there are sublinear functions $h,g$ with $g\in O(u)$ such that for $\widehat{x},\widehat{w}\in\widehat{X}$ and $x^{1}\in X_{1}$, we have $d_{Y_{1}}\big(\phi_{\widehat{x}}(x^{1}),\phi_{\widehat{w}}(x^{1})\big)\leq h\big(|\widehat{x}|\vee|\widehat{w}|)\vee g(|x^{1}|)$. We show it for $\widehat{w}=\widehat{0}$, which by the triangle inequality implies the general case.
[…]
$d_{Y_{1}}\big(\phi_{\widehat{x}}(x^{1}),\phi_{\widehat{w}}(x^{1})\big)\leq d_{Y}\big(\phi(x^{1},\widehat{x}),\phi(x^{1},\widehat{w})\big)\leq Ld_{\widehat{X}}(\widehat{x},\widehat{w})+u(|\widehat{x}|\vee|\widehat{w}|)+u(|x^{1}|),$
hence we can take $h$ to be bounded near $\widehat{0}$. By defining $\tilde{h}(r):\sup_{x\leq r}=h(r)$ we may assume $h$ is nondecreasing, and by Lemma 2.2 we conclude that one can take $h$ to be subadditive and nondecreasing. We let $v=h+g$, finishing the proof. ∎
2.D. Proof of Theorem A
We recall the three definitions below from [KKL98].
Definition 2.19.
Definition 2.20.
Definition 2.21.
A space $X$ is of coarse type I (resp. II) if it is geodesic and every asymptotic cone of $X$ is of type I (resp. II).
Symmetric spaces of noncompact type and high rank are coarse type II [KL97], and so are Euclidean buildings of high rank and cocompact affine Weyl group. Examples of spaces of coarse type I are Gromov hyperbolic groups whose Gromov boundary contains at least 3 points. These examples are enough for all our purposes in this paper, which arise in Section 3 in the proofs of Corollary B and Theorem 3.16.
In order to include Euclidean factors (or, more generally, factors with asymptotic cones homeomorphic to Euclidean space) we will need a slight variation of Theorem C.
@@ 2.D. Proof of Theorem A
Let $X:=\overline{X}\times Z,\ Y:=\overline{Y}\times W$ be two products of geodesic metric spaces, and let $\phi:X\rightarrow Y$ be an $(L,u)$-bilipschitz equivalence. Assume that in any pair of $u$-admissible cones $X_{\omega},Y_{\omega}$, there is a homeomorphism $\overline{\Phi}:\overline{X}_{\omega}\rightarrow\overline{Y}_{\omega}$ such that for the induced cone map $\Phi:=\operatorname{Cone}(\phi):X_{\omega}\rightarrow Y_{\omega}$, the following diagram commutes:
$X_{\omega}$$Y_{\omega}$$\overline{X}_{\omega}$$\overline{Y}_{\omega}$$\scriptstyle{\Phi}$$\scriptstyle{\overline{\pi}_{\omega}}$$\scriptstyle{\overline{\pi}_{\omega}}$$\scriptstyle{\overline{\Phi}}$
@@
Then there is an $O(u)$-bilipschitz equivalence $\overline{\phi}:\overline{X}\rightarrow\overline{Y}$ such that the diagram
$X¯{\overline{X}}Y¯{\overline{Y}}$$\scriptstyle{\phi}$$\scriptstyle{\overline{\pi}}$$\scriptstyle{\overline{\pi}}$$\scriptstyle{\overline{\phi}}$
commutes up to sublinear error $v$ that is nondecreasing and subadditive.
We omit the proof, which is very similar to the proof of Theorem C.
Remark 2.23.
There is one essential difference between the proof of Proposition 2.22 and that of Theorem C: in the former, the constants $L_{1}$ and $\epsilon$ in the definition of compressed and uncompressed pairs are not chosen freely. Rather, we can only guarantee that there are such constants $L_{1}>L$ and $\epsilon<L_{1}^{-1}$ for which the proof carries. The reason is that in the asymptotic cones, one does not have good enough control over the $Z$ factor. The result is that the multiplicative (Lipschitz) constant of $\overline{\phi}$ may be larger than $L$. We stress that $\overline{\phi}$ is still an $O(u)$-equivalence.
Proof of Theorem A.
We denote $\overline{X}:=\prod_{i=1}^{n}X_{i},\overline{Y}:=\prod_{j=1}^{m}Y_{j}$. The map $\phi:X\rightarrow Y$ is an $O(u)$-equivalence, therefore in any $u$-admissible cone $X_{\omega}$, the map $\phi_{\omega}$ is an $L$-bilipschitz homeomorphism (Lemma 2.8). The assumption on the factors then allows us to use Theorem 1.4, and conclude that $p=q,n=m,$ and that the map $\overline{\phi}_{\omega}:\pi_{\overline{Y}_{\omega}}\circ\phi_{\omega{\restriction\{0_{\mathbf{R}^{p}}\}\times\overline{X}_{\omega}}}:\overline{X}_{\omega}\rightarrow\overline{Y}_{\omega}$ is a homeomorphism that preserves the product structure.
We conclude that for $X,Y$ and $\phi$, the hypotheses of Proposition 2.22 are met, and we obtain an $O(u)$-equivalence $\overline{\phi}:\overline{X}\rightarrow\overline{Y}$ such that
$X¯{\overline{X}}Y¯{\overline{Y}}$$\scriptstyle{\phi}$$\scriptstyle{\overline{\pi}}$$\scriptstyle{\overline{\pi}}$$\scriptstyle{\overline{\phi}}$
commutes up to a sublinear error $\overline{v}$ that is subadditive and nondecreasing. The triple consisting of $\overline{x},\overline{Y}$ and $\overline{\phi}:\overline{X}\rightarrow\overline{Y}$ admit the hypotheses of Theorem 1.4, hence in any $u$-admissible cones $\overline{X}_{\omega},\overline{Y_{\omega}}$, the map $\overline{\phi}_{\omega}:\overline{X}_{\omega}\rightarrow\overline{Y}_{\omega}$ preserves the product structure. Therefore the hypotheses of Theorem C are met, and we obtain for all $i\in\{1,\ldots,n\}$, $O(u)$-equivalences $\overline{\phi}_{i}:X_{i}\rightarrow Y_{i}$ such that the diagrams
$\overline{X}=\prod_{i=1}^{n}X_{i}$$\prod_{i=1}^{n}Y_{i}=\overline{Y}$$X_{i}$$Y_{\sigma(i)}$$\scriptstyle{\phi}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{\sigma(i)}}$$\scriptstyle{\phi_{i}}$
commute up to sublinear error $v_{i}$ that is subadditive and nondecreasing. Finally, the exterior square in the diagram
$X¯{\overline{X}}Y¯{\overline{Y}}$$X_{i}$$Y_{i}$$\scriptstyle{\phi}$$\scriptstyle{\overline{\pi}}$$\scriptstyle{\overline{\pi}}$$\scriptstyle{\overline{\phi}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\overline{\phi}_{i}}$
commutes up to a sublinear error $v^{\prime}=\overline{v}+\max_{i}v_{i}$, as can be checked as follows: for every $x\in X$,
$d(\pi_{i}\circ\overline{\pi}\circ\phi(x),\overline{\phi}_{i}\circ\pi_{i}\circ\overline{\pi}(x))$ $\leqslant d(\pi_{i}\circ\overline{\pi}\circ\phi(x),\pi_{i}\circ\overline{\phi}\circ\overline{\pi}(x))$
$+d(\pi_{i}\circ\overline{\phi}\circ\overline{\pi}(x),\overline{\phi}_{i}\circ\pi_{i}\circ\overline{\pi}(x))$
$\leqslant d(\overline{\pi}\circ\phi(x),\overline{\phi}\circ\overline{\pi}(x))+v_{i}(\overline{\pi}(x))$
$\leqslant\overline{v}(|x|)+\max_{i\in\{1,\ldots,n\}}v_{i}(|x|).$
where we used the triangle inequality and then the fact that the projections $\overline{\pi}$ and $\pi_{i}$ are distance-decreasing. [^2] The real point used here is that the projections $\overline{\pi}$ and $\overline{\pi}_{i}$ are $o(r)$-lipschitz and thus the diagram commutes in the $o(r)$-lipschitz category defined by Cornulier in [COR19].
3. Completely solvable groups, Corollary B and Theorem D
3.A. Some preliminaries on Cornulier’s reduction
@@
    $H$ is an $O(\log)$-Lipschitz retract of $G$, more precisely: $\pi:G\to H$ is $O(\log)$-Lipschitz.
@@
We prove below that the groups in Example 1.11 are not quasiisometric. The proof is less involved than that of Theorem D but the main idea already intervenes, so we give it before.
@@
Let $\alpha\in(0,1)$. The groups $G^{0}_{4,9}$ and $\mathbf{R}\times G_{3,5}^{\alpha}$ are not quasiisometric, for any $\alpha\neq 1$.
@@
We first check that $\mathbf{R}\times G_{3,3}$ is the group in the class $(\mathcal{C}_{1})$ associated to $G_{4,9}^{0}$ by [COR11], and so there is a $O(\log)$-bilipschitz equivalence $\phi\colon G_{4,9}^{0}\to\mathbf{R}\times G_{3,3}$. The Lie algebra $\mathfrak{g}=\mathfrak{g}_{4,9}^{0}$ of $G_{4,9}^{0}$ is a semidirect product $\mathfrak{heis}\rtimes\delta$ where $\mathfrak{heis}$ has a basis $(e_{1},e_{2},e_{3})$ with $e_{1},e_{2}]=e_{3}$ and $\delta e_{i}=e_{i}$ if $i=1,3$ or 0 if $i=2$. The derived subalgebra is $\mathfrak{u}=\langle e_{1},e_{3}\rangle$ and $\mathfrak{g},\mathfrak{u}]=\mathfrak{u}$, so that $\mathfrak{u}$ is the Lie algebra of the exponential radical. The matrices of $\operatorname{ad}_{e_{3}}$ and $\operatorname{ad}_{e_{4}}$ in the basis $(e_{1},e_{2})$ are, respectively,
@@
From the definition of $\rho_{1}$ it follows that, using the same basis and changing the brackets, $\rho_{1}(\mathfrak{g}_{4,9})^{0}$ is the Lie algebra $\mathfrak{u}\rtimes\langle e_{4}\rangle\times\mathbf{R}\simeq\mathfrak{g}_{3,3}\times\mathbf{R}$.
Assume now towards contradiction that $G^{0}_{4,9}$ and $\mathbf{R}\times G_{3,5}^{\alpha}$ are quasiisometric; then by Cornulier’s Theorem 3.7 there is a $O(\log)$-bilispchitz equivalence $\psi\colon\mathbf{R}\times G_{3,3}\to\mathbf{R}\times G_{3,5}^{\alpha}$. By Theorem A there is a sublinear bilipschitz equivalence $\underline{\psi}\colon G_{3,3}\to G_{3,5}^{\alpha}$. However, by [PAL20b], $O(u)$-bilipschitz equivalences between Heintze groups preserve the conformal dimensions of their Gromov boundaries. $G_{3,3}$ and $G_{3,5}^{\alpha}$ are Heintze groups, and the conformal dimensions are
@@
3.C. Proof of Corollary B and Theorem D